THE VISIBLE SPECTRUM OF IRIDIUM MONOHYDRIDE AND MONODEUTERIDE.

A. G. ADAM, and \underline{A. D. GRANGER}, Chemistry Department, and Centre for Lasers, and Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, NB, E3B 5A3; and C. LINTON, and D. W. TOKARYK, Physics Department, and Centre for Lasers, and Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, NB, E3B 5A3.

Iridium-containing moieties are prevalent in asymmetric catalytic systems. For this reason, diatomic Ir containing species have been the subject of both theoretical and experimental studies. We have investigated IrH/D using our laser ablation molecular jet apparatus. After searching from 380 - 820 nm, eight rotationally-resolved bands for both IrH and IrD were observed. The spectra were determined to belong to several ^{3}Φ_{4} - X^{3}Φ_{4} transitions. No other spin-orbit components were observed in either the excited or ground states. A global fit of the bands was performed for each molecule. The molecular parameters have been compared with those of the valence isoelectronic molecules, CoH and RhH. Ground state vibrational frequencies have been determined, from dispersed fluorescence spectra, to be ΔG_{1/2} = 2140(11) cm^{−1} for IrH and ω_{e}′′ = 1609(2) cm^{−1} and ω_{e}x_{e}′′ = 19.3(5) cm^{−1} for IrD. The ground state bond length of IrH was determined to be 1.603 Å somewhat in agreement with the theoretical prediction of r_{e}′′ = 1.565 Å of Dai and Balasubramanian^{a}.

aD. Dai and K. Balasubramanian New. J. Chem. \underline{15}, 721-726 (1991)